Jermaine wants to know how fast he is traveling in a school bus, but he cannot see the speedometer. He records that the bus travels 1 mile in 75 seconds. How fast is Jermaine traveling in miles per hour?
Found 2 solutions by rothauserc, Boreal:
Answer by rothauserc(4718)

rate * time = distance
:
we are given time = 75 seconds and distance = 1 mile
:
we want the rate to be expressed as miles per hour so we convert 75 seconds to a fraction of an hour
:
there are 60 * 60 = 3600 seconds in an hour, therefore
:
75 / 3600 = 0.020833333 or approximately 0.0208 hours
:
**************************************************************
rate = 1 / 0.0208 = 48.0769 or approximately 48 miles per hour
**************************************************************
:

Answer by Boreal(15235)

1 mile is 75 seconds is x miles/3600 seconds, an hour
1/75=x/3600
3600=75x
x=48 mph.